Ask your own question, for FREE!
Mathematics 16 Online
OpenStudy (anonymous):

If the probability of seeing moose in a day in a certain region of Alaska is 0.29, what is the probability of not seeing any moose in a day there? My Ans is -0.29 is that correct

OpenStudy (amistre64):

1 - .29 id assume

OpenStudy (amistre64):

the opposite of -.29 lol ..... its 'accurate' but prolly not practical

OpenStudy (anonymous):

Thanks:)

OpenStudy (jit4won):

1-0.29 =0.71

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!